jake is 5 years older than his sister three years ago jake was twice his sisters age how old is jakes siter now?
Answer by stanbon(75887)

jake is 5 years older than his sister
three years ago jake was twice his sisters age
how old is jakes siter now?
------------
Equations:
j = s + 5
j-5 = 2(s-5)
------
Substitute for "j" and solve for "s"::
s+5-5 = 2s-10
s = 2s-10
s = 10 (sister's age now)
----
Solve for "j":
j = s+5
j = 15 (jake's age now)
========================
Cheers,
Stan H.
=================